OpenStudy (anonymous):

Simplify. −1w + 4w4 + y4 – 3y4 – 2w – 3w4 A. −2w4 – 2y4 B. −w4y4 – 3w C. w4 – 3w – 2y4 D. 12w4 + 2w – 3y4

OpenStudy (michele_laino):

please you have to sum similar monomials

OpenStudy (anonymous):

(-1w+2w)+(4w4-3w4)-(y4_3y4) is this right?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

please first parenthesis is -1w-2w

OpenStudy (anonymous):

I got 1w+1w4-3y4

OpenStudy (michele_laino):

sorry , I got another answer,

OpenStudy (anonymous):

Can you tell me how you got another answer?

OpenStudy (michele_laino):

don't worry, now first term is -1w-2w, are you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok!, second term is right, namely 4w^4-3w^4

OpenStudy (michele_laino):

third term is -(3y^4-y^4) are you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! so retry please

OpenStudy (anonymous):

how do you add and subtract them together?

OpenStudy (michele_laino):

for example -1w-2w=-3w, because -1+(-2)=-3

OpenStudy (michele_laino):

literal parts is the same, you have only to add the numerical parts each other, in order to add similar monomials

OpenStudy (michele_laino):

is it clear?

OpenStudy (anonymous):

Not really

OpenStudy (michele_laino):

another example: -3x+5x=2x, because the literal part is the common literal part, namely x, whereas the coefficient is 2, because -3+5=2, so 2x

OpenStudy (anonymous):

So its between b and c

OpenStudy (michele_laino):

please let us continue as before, the first term is -3w, are you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok!, the second term is 4w^4-3w^4=w^4, are you agree?

OpenStudy (anonymous):

I thought it was 1w4

OpenStudy (michele_laino):

yes it the same thing, in general when the coefficient is 1, you can omit it

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

now the third term is -(3y^4-y^4)=-(2y^4)=-2y^4, are you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

perfect, now you have to sum these three terms: -3w; w^4 -2y^4 note that they are not similar monomials, so whe we add them each other, we can only write them one near to other, so?

OpenStudy (anonymous):

so its c

OpenStudy (michele_laino):

well done!, that's right!

OpenStudy (anonymous):

Thank you. You were alot of help

OpenStudy (michele_laino):

Thank you!